# The Statistical Fairness-Accuracy Frontier

### v1

@@ 6 Conclusion
Our analysis demonstrates that operationalizing the FA frontier with finite data poses unique challenges. Notably, departing from the population ideal introduces asymmetries across groups. These asymmetries grow with group heterogeneity, and such settings are precisely where fairness considerations are often most pressing. More specifically, we show that, when the covariance structure is known, the optimal sampling allocation prioritizes groups with greater variance or higher weight in the objective. When the covariance structure is unknown, however, a new challenge emerges: the bias term, which motivates balancing sample sizes across groups. In both settings, failing to account for these asymmetries can yield solutions that deviate meaningfully from the intended population-level trade-off.
@@ A.1 Auxiliary Results
Lemma 3 (Mourtada [2022] Corollary 4 generalized to spherical covariance).
@@
By Mourtada [2022] Corollary 4 taking $q=4$, if $n_{g}\geq 48/\alpha_{g}$,
@@
Under Assumption 4, there exists a universal constant $C>0$ such that:
@@
By standard subgaussian matrix concentration (see, e.g., Vershynin [2018] Remark 4.7.3), for all $\delta\in(0,1)$, there exists an absolute constant $C>0$ such that:
@@
$s\leq CK_{g}^{2}\left(\sqrt{\frac{d+u}{n_{g}}}+\frac{d+u}{n_{g}}\right)\left\lVert\Sigma_{g}\right\rVert\leq 2CK_{g}^{2}\max\left\{\sqrt{\frac{d+u}{n_{g}}},\frac{d+u}{n_{g}}\right\}$ (37)
@@
    If $\sqrt{\frac{d+u}{n_{g}}}\geq\frac{d+u}{n_{g}},$ (37) is satisfied if $s\leq 2CK_{g}^{2}\sqrt{\frac{d+u}{n_{g}}}$ which occurs when $u\geq\frac{s^{2}n_{g}}{4C^{2}K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}}-d$.
@@
    If $\sqrt{\frac{d+u}{n_{g}}}<\frac{d+u}{n_{g}},$ (37) is satisfied if $s\leq 2CK_{g}^{2}\sqrt{\frac{d+u}{n_{g}}}$ which occurs when $u\geq\frac{sn_{g}}{2CK_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}-d$.
@@
Define the critical point $s_{d}\coloneqq\min\{s_{1},s_{2}\}$ and split the integral
@@
$\int_{0}^{s_{d}}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq\int_{0}^{s_{d}}s^{3}ds\leq\frac{1}{4}\max\{s_{1}^{4},s_{2}^{4}\}=\frac{1}{4}\max\left\{\left(\frac{d}{C^{\prime}n_{g}}\right)^{2},\left(\frac{d}{C^{\prime}n_{g}}\right)^{4}\right\}$
For the second term, observe
$\int_{s_{d}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq\max\left\{\int_{s_{1}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds,\int_{s_{2}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\right\}.$
We then bound each integral:
$\int_{s_{1}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq\int_{s_{1}}^{\infty}s^{3}\exp\left(\frac{-C^{\prime}n_{g}s^{2}}{K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}}\right)ds=\frac{K_{g}^{8}(d+1)\exp(-d)\left\lVert\Sigma_{g}\right\rVert^{4}}{2C^{\prime 2}n_{g}^{2}}$
$\int_{s_{2}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq\int_{s_{2}}^{\infty}s^{3}\exp\left(\frac{-C^{\prime}n_{g}s}{K_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}\right)ds=\frac{K_{g}^{8}\left(d^{3}+3d^{2}+6d+6\right)\exp({-d}){\left\lVert\Sigma_{g}\right\rVert}^{4}}{C^{\prime 4}n_{g}^{4}}.$
Therefore, in either case, we conclude that, for an absolute constant $C>0$,
@@
For the cross terms, by independence, for $i\neq j$,
@@
By our subgaussianity assumption 4 and standard moment bounds for subgaussian vectors (see, e.g., Vershynin [2018]), we have, for some absolute constant $C>0$:
@@
Put the pieces together, we have
$\mathbb{E}\left[\operatorname{Tr}\left(\widehat{\Sigma}_{g}^{2}\right)\right]\leq\operatorname{Tr}\left(\Sigma_{g}^{2}\right)+\frac{CK_{g}^{4}}{n_{g}}\operatorname{Tr}(\Sigma_{g})^{2}$
as desired. $\blacksquare$
Lemma 6 (Assouad’s Lemma [Yu, 1997]).
@@ A.3.1 Proof of Theorem 1
Substituting $\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{b}$ and using the decomposition
@@
$\tilde{\beta}_{\lambda}-\beta_{\lambda}=\Sigma_{\lambda}^{-1}\left(\lambda\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\cdot\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)\Sigma_{b}\widehat{\Sigma}_{b}\cdot\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right).$
Let $A\coloneqq{\Sigma}_{\lambda}^{-1}\Sigma_{g}{\Sigma}_{\lambda}^{-1}$. By the mean-zero and independence properties of the noise, the cross terms cancel in the squared norm, which yields
@@ A.3.2 Proof of Corollary 2
Then, by whitening and applying Mourtada [2022, Theorem 3], we replace (42) with the following bound:
@@ A.3.3 Proof of Proposition 2
Next, using the variant of Woodbury identity $(A-B)^{-1}=A^{-1}+A^{-1}B(A-B)^{-1}$, we obtain
@@ A.4.1 Proof of Proposition 3
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ $\leq\lambda^{2}\frac{\sigma^{2}}{n_{r}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{r}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{b}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]$
@@
with $A=\tilde{\Sigma}_{g}^{-2}$ and $x_{i}=\tilde{X_{i}}$. This gives:
@@ A.4.2 Proof of Theorem 2
We show a slightly more general result for the upper bound. In fact, we relax the (25) to
@@
and also drop the assumption $n_{g}\geq K_{g}^{4}d$, and show the upper bound
@@
Take $\xi$ and $\xi^{\prime}$ to be neighbors that differ only in the $i$-th coordinate of $\xi_{r}$, i.e., $\xi_{r,i}=1,\xi_{r,i}=-1$ and $\xi_{r,-i}=\xi^{\prime}_{r,-i}$. Recalling well-known results of the KL divergence of multivariate Gaussians (see, e.g., Zhang et al. [2023]), conditional on observing a single observation $X_{i}$, the KL divergence between $\mathbb{P}_{\xi_{r}}=\mathcal{N}(h_{r}X_{i},\sigma^{2}),\mathbb{P}_{\xi^{\prime}_{r}}=\mathcal{N}(-h_{r}X_{i},\sigma^{2})$ is given by
@@ A.4.3 Proof of Theorem 3
Substituting into (72a) and (72b) into (71) and applying the assumptions $\left\lVert\Sigma_{g}\right\rVert\leq\mathrm{P}_{g}^{2}$ and $\lambda_{\min}\left({\Sigma_{\lambda}}\right)\geq\rho_{\lambda}^{2}$, we reach
@@
By Sherman Morrison, $\left(\Sigma_{r}-h_{r}u_{i}u_{i}^{\top}\right)^{-1}=\Sigma_{r}^{-1}\frac{h_{r}}{1-\alpha}(\Sigma_{r}^{-1}u_{i}u_{i}^{\top}\Sigma_{r}^{-1})$. Substituting and simplifying, we reach
@@
Observe $\alpha\geq 0$ since $\Sigma_{r}$ is PSD. By (79), $\Sigma_{r}^{-1}\preceq\frac{1}{0.9\rho_{r}^{2}}I_{d}$ thus $\alpha\leq\frac{h_{r}}{0.9\rho_{r}^{2}}$. By (78), it follows that $0\leq\alpha\leq\frac{1}{9d}\leq\frac{1}{9}$ for $d\geq 1$. Define $\gamma\coloneqq\frac{2\alpha}{1-\alpha}$ where $\gamma\in[0,1/4]$. Using the fact that, for $\gamma\in[0,1/4]$, $\log(1+\lambda)+\lambda\leq\frac{\lambda^{2}}{2}$ and the fact that, for $\alpha\in(0,1/9)$, $\frac{\alpha^{2}}{(1-\alpha)^{2}}\leq\frac{81}{64}\alpha^{2}$,
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})=\frac{1}{2}\left[-\log(1+\lambda)+\lambda\right]\leq\frac{\lambda^{2}}{4}=\frac{\alpha^{2}}{(1-\alpha)^{2}}\leq\frac{81}{64}\alpha^{2}\leq\frac{81}{64}h_{r}^{2}\left\lVert\Sigma_{r}^{-1}\right\rVert_{2}^{2}\leq\frac{25}{16}\frac{h_{r}^{2}}{\rho_{r}^{4}}$
@@
By 1, over $n_{r}$ independent samples from group $r$,
@@
$\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}=\left({A^{(\xi)}}^{-1}-{A^{(\xi^{\prime})}}^{-1}\right)(1-\lambda)\Sigma_{b}(\beta_{b}-\beta_{r})\Sigma_{b}(\beta_{b}-\beta_{r}).$
@@ A.4.4 Proof of Proposition 4
$\mathbb{E}\left[\widehat{\beta}_{\lambda}-\beta_{\lambda}\right]=\mathbb{E}\left[\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})+(1-\lambda)\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda})\right)\right]$ (88)
Woodbury matrix identity, we can write
@@
Next, we bound the two terms on the right hand side separately. First, notice that

### v2

@@ 8 Conclusion
Our analysis demonstrates that operationalizing the FA frontier with finite data poses unique challenges. Notably, departing from the population ideal introduces asymmetries across groups. These asymmetries grow with group heterogeneity, and such settings are precisely those in which fairness considerations are often most pressing. More specifically, we show that, when the covariance structure is known, the optimal sampling allocation prioritizes groups with greater variance or higher weight in the objective. When the covariance structure is unknown, however, a new challenge emerges: the bias term, which motivates balancing sample sizes across groups. In both settings, failing to account for these asymmetries can yield solutions that deviate meaningfully from the intended population-level trade-off.
@@ A.1 Auxiliary Results
Lemma 3 (mourtada2022exact Corollary 4 generalized to spherical covariance).
@@
By mourtada2022exact Corollary 4 taking $q=4$, if $n_{g}\geq 48/\alpha_{g}$,
@@
Suppose Assumption 4 holds and $n_{g}\geq d$. Then, there exists a universal constant $C>0$ such that:
@@
By standard subgaussian matrix concentration [see, e.g., vershynin2018high, Remark 4.7.3], for all $\delta\in(0,1)$, there exists an absolute constant $C>0$ such that:
@@
$s\leq CK_{g}^{2}\left(\sqrt{\frac{d+u}{n_{g}}}+\frac{d+u}{n_{g}}\right)\left\lVert\Sigma_{g}\right\rVert\leq 2CK_{g}^{2}\max\left\{\sqrt{\frac{d+u}{n_{g}}},\frac{d+u}{n_{g}}\right\}\left\lVert\Sigma_{g}\right\rVert$ (38)
@@
    If $\sqrt{\frac{d+u}{n_{g}}}\geq\frac{d+u}{n_{g}},$ (38) is satisfied if $s\leq 2CK_{g}^{2}\sqrt{\frac{d+u}{n_{g}}}\left\lVert\Sigma_{g}\right\rVert$ which occurs when $u\geq\frac{s^{2}n_{g}}{4C^{2}K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}}-d$.
@@
    If $\sqrt{\frac{d+u}{n_{g}}}<\frac{d+u}{n_{g}},$ (38) is satisfied if $s\leq 2CK_{g}^{2}{\frac{d+u}{n_{g}}}\left\lVert\Sigma_{g}\right\rVert$ which occurs when $u\geq\frac{sn_{g}}{2CK_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}-d$.
@@
Define the critical point $s_{d}\coloneqq\max\{s_{1},s_{2}\}$ and split the integral
@@
$\int_{0}^{s_{d}}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq\int_{0}^{s_{d}}s^{3}ds\leq\frac{1}{4}\max\{s_{1}^{4},s_{2}^{4}\}=\frac{K_{g}^{8}\left\lVert\Sigma_{g}\right\rVert^{4}}{4}\max\left\{\left(\frac{d}{C^{\prime}n_{g}}\right)^{2},\left(\frac{d}{C^{\prime}n_{g}}\right)^{4}\right\}$
For the second term, we consider two cases of $s_{d}=s_{1}$ and $s_{d}=s_{2}$:
$\int_{s_{1}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq 2\int_{s_{1}}^{\infty}s^{3}\exp\left(d-\frac{C^{\prime}n_{g}s^{2}}{K_{g}^{4}\left\lVert\Sigma_{g}\right\rVert^{2}}\right)ds=\frac{K_{g}^{8}(d+1)\left\lVert\Sigma_{g}\right\rVert^{4}}{C^{\prime 2}n_{g}^{2}}$
$\int_{s_{2}}^{\infty}\mathbb{P}\left(\left\lVert\widehat{\Sigma}_{g}-\Sigma_{g}\right\rVert\geq s\right)s^{3}ds\leq 2\int_{s_{2}}^{\infty}s^{3}\exp\left(d-\frac{C^{\prime}n_{g}s}{K_{g}^{2}\left\lVert\Sigma_{g}\right\rVert}\right)ds=\frac{2K_{g}^{8}\left(d^{3}+3d^{2}+6d+6\right){\left\lVert\Sigma_{g}\right\rVert}^{4}}{C^{\prime 4}n_{g}^{4}}.$
Therefore, in either case, using $n_{g}\leq d$, we conclude that, for an absolute constant $C>0$,
@@
For the cross-terms, by independence, for $i\neq j$,
@@
By our subgaussianity assumption (Assumption 4) and standard moment bounds for subgaussian vectors [see, e.g., vershynin2018high], we have, for some absolute constant $C>0$:
@@
Putting these results together yields the claim. $\blacksquare$
Lemma 6 (Assouad’s Lemma ).
@@ A.3.1 Proof of Theorem 1
Substituting $\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}$ and using the decomposition
@@
$\tilde{\beta}_{\lambda}-\beta_{\lambda}=\Sigma_{\lambda}^{-1}\left(\lambda\Sigma_{r}\widehat{\Sigma}_{r}^{-1}\cdot\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}X_{i}^{(r)}\varepsilon_{i}^{(r)}+(1-\lambda)\Sigma_{b}\widehat{\Sigma}_{b}^{-1}\cdot\frac{1}{n_{b}}\sum_{i=1}^{n_{b}}X_{i}^{(b)}\varepsilon_{i}^{(b)}\right).$
Let $A\coloneqq{\Sigma}_{\lambda}^{-1}\Sigma_{g}{\Sigma}_{\lambda}^{-1}$. By the mean-zero and independence properties of the noise, the cross-terms cancel in the squared norm, which yields
@@ A.3.2 Proof of Corollary 2
Then, by whitening and applying mourtada2022exact, we replace (43) with the following bound:
@@ A.3.3 Proof of Proposition 2
Next, using a variant of the Woodbury identity, $(A-B)^{-1}=A^{-1}+A^{-1}B(A-B)^{-1}$, we obtain
@@ A.4.1 Proof of Proposition 3
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]$ $=\lambda^{2}\frac{\sigma^{2}}{n_{r}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{r}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]+(1-\lambda)^{2}\frac{\sigma^{2}}{n_{b}^{2}}\mathbb{E}\left[\sum_{i=1}^{n_{b}}\left\lVert\tilde{X_{i}}\right\rVert_{\tilde{\Sigma}_{g}^{-2}}^{2}\right]$
@@
with $A=\tilde{\Sigma}_{g}^{-1}$ and $x_{i}=\tilde{X_{i}}$. This gives:
@@ A.4.2 Proof of Theorem 2
We show a slightly more general result for the upper bound. In fact, we relax the condition (26) to
@@
and also drop the assumption $n_{g}\geq K_{g}^{4}d$, and establish the upper bound
@@
To see this, take $\xi$ and $\xi^{\prime}$ to be neighbors that differ only in the $i$-th coordinate of $\xi_{r}$, i.e., $\xi_{r,i}=1,\xi_{r,i}=-1$ and $\xi_{r,-i}=\xi^{\prime}_{r,-i}$. Recalling well-known results for the KL divergence of multivariate Gaussians [see, e.g., zhang2023properties], conditional on observing a single observation $X_{i}$, the KL divergence between $\mathbb{P}_{\xi_{r}}=\mathcal{N}(h_{r}X_{i},\sigma^{2}),\mathbb{P}_{\xi^{\prime}_{r}}=\mathcal{N}(-h_{r}X_{i},\sigma^{2})$ is given by
@@ A.4.3 Proof of Theorem 3
Substituting (73a) and (73b) into (72) and applying the assumptions $\left\lVert\Sigma_{g}\right\rVert\leq\mathrm{P}_{g}^{2}$ and $\lambda_{\min}\left({\Sigma_{\lambda}}\right)\geq\rho_{\lambda}^{2}$, we reach
@@
By the Sherman-Morrison lemma, $\left(\Sigma_{r}-h_{r}u_{i}u_{i}^{\top}\right)^{-1}=\Sigma_{r}^{-1}\frac{h_{r}}{1-\alpha}(\Sigma_{r}^{-1}u_{i}u_{i}^{\top}\Sigma_{r}^{-1})$. Substituting and simplifying, we reach
@@
Observe $\alpha\geq 0$ since $\Sigma_{r}$ is PSD. By (80), $\Sigma_{r}^{-1}\preceq\frac{1}{0.9\rho_{r}^{2}}I_{d}$ thus $\alpha\leq\frac{h_{r}}{0.9\rho_{r}^{2}}$. By (79), it follows that $0\leq\alpha\leq\frac{1}{9d}\leq\frac{1}{9}$ for $d\geq 1$. Define $\gamma\coloneqq\frac{2\alpha}{1-\alpha}$ where $\gamma\in[0,1/4]$. Using the fact that, for $\gamma\in[0,1/4]$, $\log(1+\gamma)+\gamma\leq\frac{\gamma^{2}}{2}$ and the fact that, for $\alpha\in(0,1/9)$, $\frac{\alpha^{2}}{(1-\alpha)^{2}}\leq\frac{81}{64}\alpha^{2}$,
$D_{\text{KL}}(\mathbb{P}^{(\xi)}\;\|\;\mathbb{P}^{(\xi^{\prime})})=\frac{1}{2}\left[-\log(1+\gamma)+\gamma\right]\leq\frac{\gamma^{2}}{4}=\frac{\alpha^{2}}{(1-\alpha)^{2}}\leq\frac{81}{64}\alpha^{2}\leq\frac{81}{64}h_{r}^{2}\left\lVert\Sigma_{r}^{-1}\right\rVert_{2}^{2}\leq\frac{25}{16}\frac{h_{r}^{2}}{\rho_{r}^{4}}$
@@
By 2, over $n_{r}$ independent samples from group $r$,
@@
$\beta_{\lambda}^{(\xi)}-\beta_{\lambda}^{(\xi^{\prime})}=\left({A^{(\xi)}}^{-1}-{A^{(\xi^{\prime})}}^{-1}\right)(1-\lambda)\Sigma_{b}(\beta_{b}-\beta_{r}).$
@@ A.4.4 Proof of Proposition 4
$\mathbb{E}\left[\widehat{\beta}_{\lambda}-\beta_{\lambda}\right]=\mathbb{E}\left[\widehat{\Sigma}_{\lambda}^{-1}\left(\lambda\left(\widehat{\Sigma}_{r}-\Sigma_{r}\right)(\beta_{r}-\beta_{\lambda})+(1-\lambda)\left(\widehat{\Sigma}_{b}-\Sigma_{b}\right)(\beta_{b}-\beta_{\lambda})\right)\right].$ (89)
By the Woodbury matrix identity, we can write
@@
Next, we bound the two terms on the right-hand side separately. First, notice that